### v1

@@ 5 Conclusion
In this paper, we present our methods for the 7th ABAW Multi-task Learning Challenge (MTL) and the Compound Expression (CE) Recognition task. Based on our re-collected large-scale facial dataset, we train a robust facial feature extractor. Meanwhile, based on the pretrained feature extractor, we finetune it on downstream task datasets thus obtaining the high-quality facial expression feature. For the MTL task, we devise the feature fusion model to fuse the features extracted from other affective analysis tasks with features extracted from the current model, which improves the model performance for EXPR, Valence, and Arousal estimation models. Moreover, our temporal convergence model captures the expression dynamic changes over sequence frames, enhancing the model performance for Valence and Arousal estimation models. For the CE task, we introduce the curriculum learning strategy that allows the model to effectively transfer its learned features and adapt to the more complex task of identifying blended or compound expressions. In addition, we employ different data augmentation techniques to generate new compound expressions, maximizing the utility of all available data.

### v3

@@ 5 Conclusion
In this paper, we present our methods for the 7th ABAW Multi-task Learning Challenge. Based on our re-collected large-scale facial dataset, we train a robust facial feature extractor. Meanwhile, based on the pretrained feature extractor, we finetune it on downstream task datasets thus obtaining the high-quality facial expression feature. We devise the feature fusion model to fuse the features extracted from other affective analysis tasks with features extracted from the current model, which improves the model performance for EXPR, Valence, and Arousal estimation models. Moreover, our temporal convergence model captures the expression dynamic changes over sequence frames, enhancing the model performance for Valence and Arousal estimation models. Comprehensive experiments demonstrate the superiority of our method compared to other solutions.
Acknowledgements. This research is funded in part by the National Key R&D Program of China (No. 2022YFF09022303 to Lincheng Li), ARC-Discovery grant (DP220100800 to XY), and ARC-DECRAgrant (DE230100477 to XY). The first author is funded by the CSC scholarship and Data61-topup scholarship.